OpenStudy (anonymous):

Please help.. Johnny has been planting blueberries in a garden in his backyard. Johnny started with blueberries in one square foot of garden modeled by the function s(x) = 10. He is allowed to convert more of the garden, and each additional square foot he maintains allows his blueberry plants to produce more by a rate of a(x) = 1.4x-2. Explain to Johnny how to create an equation to predict the number of blueberries he can expect based on the number of square feet he maintains. Describe how to determine the number of blueberries he will grow with 5 square feet. Use complete sentences.

OpenStudy (anonymous):

@Preetha @Hero @iambatman

OpenStudy (anonymous):

@abb0t @cwrw238

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

Can you please help? c:

undeadknight26 (undeadknight26):

Sorry mate i am a little busy at the moment D:

OpenStudy (anonymous):

Oh okay :c Do you know anyone who could?

undeadknight26 (undeadknight26):

@dan815 @Hero These are 2 very smart people :D

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Think you can help again?

OpenStudy (dan815):

u didnt tag me this time =[

OpenStudy (dan815):

was it my face??

OpenStudy (anonymous):

I'm sorry, you helped me on the last two so I figured I bothered you enough :c

OpenStudy (dan815):

haha im jk

OpenStudy (anonymous):

If you don't mind helping me on this one that'd be awesome ^.^

OpenStudy (dan815):

Please help.. Johnny has been planting blueberries in a garden in his backyard. Johnny started with blueberries in one square foot of garden modeled by the function s(x) = 10. He is allowed to convert more of the garden, and each additional square foot he maintains allows his blueberry plants to produce more by a rate of a(x) = 1.4x-2. Explain to Johnny how to create an equation to predict the number of blueberries he can expect based on the number of square feet he maintains. Describe how to determine the number of blueberries he will grow with 5 square feet. Use complete sentences.

OpenStudy (dan815):

lem me go put my mango downstairs

OpenStudy (anonymous):

Okay .-.

OpenStudy (dan815):

okay back

OpenStudy (anonymous):

Did your mango have fun?

OpenStudy (dan815):

s(x) = 10 a(x) = 1.4x-2 <---- rate by which a blue berry production increases per plant

OpenStudy (dan815):

a(x) = 1.4x-2 <---- rate by which a blue berry production increases per sqr foot*

OpenStudy (anonymous):

okay. So what do we do with s(x)

OpenStudy (dan815):

okay so since a(x) is the rate Total production, for x>=1 s(x) * a(x) = Total production

OpenStudy (dan815):

err wait no

OpenStudy (anonymous):

*waits* .-.

OpenStudy (dan815):

s(x) + s(x) * a(x) = Total production ^--- starting with this no matter what increasing by a(x) factor on what he can produce at first so s(x) * a(x)

OpenStudy (anonymous):

so 20 * a(x)= total

OpenStudy (dan815):

Let P(x) be the production function per square root for x>=1 P(x)=s(x)+s(x)*a(x) = 10 + 10*(1.4x-2)

OpenStudy (dan815):

no do PEDMAS

OpenStudy (anonymous):

But do we know what x is?

OpenStudy (dan815):

x is the number of square feet you want to maintain

OpenStudy (anonymous):

so 5

OpenStudy (dan815):

no x is the additionat square feet you want to maintain

OpenStudy (dan815):

so when it asks how much he will maintain for 5, that means he's maintaining 4 more squre feet so

OpenStudy (dan815):

10+10(1.4(4)-2)=Production

OpenStudy (anonymous):

p=23.6

OpenStudy (anonymous):

wait no 72 sorry

OpenStudy (anonymous):

So he can expect 72 blueberries?

OpenStudy (dan815):

do the arithmetic right

OpenStudy (anonymous):

46?

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Hero Can you please help me finish this

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

:-|